# FTPI-MINN-25-11

### v1

@@ 2.2.2 Anomaly in the superpotential
This anomaly in ${\mathcal{N}}=(1,1)$ theories was first discovered in [30] in a model with a single real boson field (and its fermion superpartner). Here we will outline this derivation assuming that the number of fields is arbitrary. We start from the Lagrangian (2.1) implying the classical expression for the supercurrent quoted in (2.2). Classically,
@@
due to the fact that $\gamma_{\mu}\gamma^{\nu}\gamma^{\mu}=0$ in two dimensions. As we will see, the right-hand side of (2.15) changes upon the UV regularization. The simplest one is the dimensional regularization to be used below in which $\gamma_{\mu}\gamma^{\nu}\gamma^{\mu}=D-2$. (Note also that in [30] a regularization through higher derivatives provided an additional verification.)
To carry out the one-loop calculation we apply the background field technique; namely we substitute the field $\phi^{a}$ by its background and quantum parts, $\phi^{a}$ and $\chi^{a}$, respectively, [^2] Why one loop is sufficient will be explained below.
@@ 3.2 $\mathcal{N}=(1,1)$ supersymmetrization
Figure 1: (Super)Potentials for the MSTB model for $\lambda=m=1$. (1(a)) Potential (3.2) at $\kappa=0.8$. (1(b)) and 1(c) Superpotential (3.8) at $\kappa=0.3$ (the same plot from two different points of view). The values of $\kappa$ are chosen so that the characteristic features are most prominent visually. Trajectories of the degenerate kinks (3.15) are shown by solid lines, magenta and green respectively. Trajectory of the sphaleron (3.32) is shown as a dashed red line. On (1(b)) and (1(c)), the BPS (and semi-BPS) trajectories can be seen as trajectories of viscous honey trickling down on the superpotential landscape (different trajectories correspond to different infinitesimal speeds).
@@ 3.5.2 Scalar VEVs and the fate of the ground states
Now let us consider some implications of the correction to the superpotential, Eq. (3.47). We will focus on the case $|\kappa|\leqslant 1$ relevant for the study of degenerate kinks. As in Eq. (2.26), we introduce full superpotential:
@@
The second term here is the result of the quantum correction (3.47). Note that it is strictly positive, and for large enough $\lambda$ this equation might not have real solutions.
@@
$\lambda_{\text{crit,2}}=\frac{m}{3}\sqrt{\frac{\pi}{2}\left(2+\kappa^{2}+\kappa\sqrt{3(4-\kappa^{2})}\right)}\,$, (3.56)
@@ 3.6 Instanton (bosonic sector)
Now it is time to discuss the instanton configuration that interpolates between the two degenerate kinks discussed above. This study was initiated in [17]. For now we will focus only on the bosonic sector; fermionic sector will be discussed later.
@@
Physically, one way to think about an instanton is to consider it a classical trajectory (or a field profile) that interpolates between different ground states of the theory. At the same time, in the model at hand we wish to consider an instanton that interpolates between two kinks. To consolidate these notions, we can put our theory in boundary conditions with a twist spatial with respect to the $\mathbb{Z}_{2}$ transformation $S:\ \phi_{1}\to-\phi_{1}$, see Eq. (3.3a). Recall that this $S$ is a valid symmetry of the bosonic sector of the theory; the fermionic sector will be discussed below. Such a twist puts the theory in the non-trivial topological sector; in particular, the ground states of this theory necessarily have non-trivial topology as well. Since the two degenerate kinks are the global minima of the action with the given boundary conditions (they saturate the Bogomolny bound), we conclude that, in fact, these two kinks are the ground states in this sector.
@@
3.6.1 Bosonic worldline effective action in the limit $\kappa\to 1$
@@
With the ansatz (3.64), it is straightforward to obtain the effective action for the $\beta(\tau)$. To this end we plug the ansatz (3.64) into the Euclidean action (3.63), expand in the powers of $\beta$, and keep only a few lowest order relevant terms. Recall that we assume $|\beta(\tau)|<\alpha\ll 1$. Integrating over $z$, we obtain [^8] See [29] for details. the effective quantum mechanical action for $\beta(\tau)$ (Euclidean notation):
$S_{\beta}^{\text{eff}}=\frac{m}{\lambda^{2}}\frac{1}{2}\dot{\beta}^{2}+V^{\text{eff}}(\beta)\,$, (3.65)
@@
Therefore, when $\alpha$ is not small, interference effects can become noticeable. In particular, the may be a certain value of the parameter $\alpha$, at which two different one-instanton amplitudes cancel because of the non-trivial relative phase.
@@ 3.6.2 Instanton in the limit $\kappa\to 0$
At the point $\kappa=0$, the kinks disappear, but the instanton solution is formally there — it becomes the global vortex of the Abelian Higgs model. Its action diverges logarithmically in this limit,
@@ 3.7.1 Overall translational modes of the kink-instanton system
We expect that the kink zero modes from Sec. 3.3.3 smoothly continue through the instanton. This is an overall fermionic mode of the configuration from Fig. 7. When we switch on instantons, both in plain and topologically nontrivial backgrounds, we must perform the Wick rotation, i.e. pass to Euclidean time. Then one fermion zero mode on the kink we discussed becomes a zero mode on the domain line in the plane $\{\tau,z\}$ ($\tau$ is the Euclidean time). Its profile function depends on $z$ but is $t$-independent. Instanton deforms this line locally, near its center $z_{0}$ and $\tau_{0}$, but the boundary behavior of the background field remains the same, which means that the profile is deformed around $z_{0},\tau_{0}$ but the mode survives.
@@ 3.7.2 Modes localized on the instanton
Unfortunately, index theorems cannot be directly applied in the present case, because one mass eigenvalue vanishes at a line in the field space, see footnote 4 on p. 4. This does not allow us to rigorously prove our statement (unlike the $\mathcal{N}=2$ case treated in Sec. 4.5.2). Still, below in Sec. 3.7.4 we present an argument, which, although conjectural, makes a plausible case. But before doing that, we need to recall some facts about $\mathcal{N}=2$ 2d Wess-Zumino models with holomorphic superpotentials.
@@ 4.1 Setup
Here, the prime represents a derivative w.r.t. $\Phi$.
@@
Superpotential (4.1) is motivated by $\mathcal{N}=1$ SQCD in 4D, with gauge group $SU(2)$ and one quark flavor. There, $\Phi$ plays the role of the meson superfield. In that context, $b=m$ is the quark mass, and $a=\Lambda^{5}$ in terms of the strong coupling scale. Generalization to the gauge group $SU(N)$ yields the superpotential
@@ 4.2 Two degenerate kinks
$\partial_{z}\theta=-2b\sin\theta\,.$ (4.12)
@@
$\theta=\pm 2\;\mathrm{arctan}(e^{-2b(z-z_{0})})\,$, (4.13)
@@
$\Phi_{\text{u-kink}}(z)=\Phi_{+}\cdot\frac{e^{2bz}+i}{e^{2bz}-i}\,$, (4.14a)
$\Phi_{\text{d-kink}}(z)=\Phi_{+}\cdot\frac{e^{2bz}-i}{e^{2bz}+i}\,.$ (4.14b)
Both kinks interpolate from $\Phi_{-}$ at $z\to$-$\infty$ to $\Phi_{+}$ at $z\to+\infty$, see Fig. 9. The first kink $\Phi=\Phi_{\text{u-kink}}(z)$ goes through the upper half-plane on the field space; the phase of $\Phi$ (i.e. $\theta$) interpolates from $\pi$ to 0 clockwise. The second kink $\Phi=\Phi_{\text{d-kink}}(z)$ goes through the lower half-plane on the field space; the phase of $\Phi$ interpolates from $\pi$ to 0 going counter-clockwise.
@@ 4.3.1 First order equations for sphaleron profile
$z-z_{0}=\frac{1}{b}\left[\Phi_{\text{sph}}(z)-\Phi_{+}\,\text{atanh}\frac{\Phi_{\text{sph}}(z)}{\Phi_{+}}\right]\,$, (4.15)
where the vacuum value $\Phi_{+}$ is defined in (4.4), and $z_{0}$ is an arbitrary constant (the sphaleron center). For a visual representation, see Fig. 9.
@@
$\Phi_{\text{sph}}(z)\approx(3az)^{1/3}\,,\quad\Phi_{\text{sph}}^{\prime}(z)\sim\frac{1}{z^{2/3}}\,.$ (4.16)
@@ 4.3.2 Sphaleron energy and BPS bound
4.3.3 Negative mode
Let us write down the second order Euler-Lagrange equation for the energy functional (4.8):
$\partial_{z}^{2}\Phi=-\partialderivative{\bar{\Phi}}V(\Phi,\bar{\Phi})=\frac{2a}{\bar{\Phi}^{3}}\left(b-\frac{a}{\Phi^{2}}\right)\,.$ (4.19)
It is easy to see that the sphaleron trajectory considered here has a negative bosonic mode. To this end we write $\Phi=\Phi_{\text{sph}}(z)+\delta\phi_{1}(z)+i\delta\phi_{2}(z)$ and plug this into the second-order equations (4.19). Near the sphaleron singularity we obtain
$\delta\phi_{2}^{\prime\prime}\approx\frac{2}{\Phi_{\text{sph}}^{6}}\delta\phi_{2}\,$, (4.20)
which evidently corresponds to negative-frequency and exponential runaway behavior.
@@ 4.3.3 Negative mode
$\phi(z)_{\text{transl}}\sim\partial_{z}\Phi_{\text{sph}}(z)\,.$ (4.21)
@@ 4.4.1 Topological charge
Here, the terms $S_{\text{kink}}^{(1,2)}$ come from $2\int dz\,d\tau\,[\partial_{\zeta}\mathcal{W}+c.c.]$; they correspond to the central charges of the static kinks between which the instanton interpolates [^12] Formally these terms are divergent, as we need to integrate the kink central charges over all of time. However, with given boundary conditions at $z\to\pm\infty$, a kink represents a ground state, so we subtract that contribution away, as in Sec. 3.. The term $S_{\text{inst}}$ gives the instanton action; it is given by
@@ 4.4.2 BPS profile equations and the singularity
Let us introduce polar coordinates on the Euclidean plane $(z,t)$ according to
$z=r\cos\alpha\,,\quad t=r\sin\alpha\,.$ (4.33)
@@
Previously we saw that the sphaleron had a singularity, which led to a divergence of the sphaleron’s energy. One might worry, what is the fate of the instanton, since the instanton configuration inevitably passes through the point $\Phi=0$ in the field space; let’s say that this happens at the origin of $(z,t)$ plane. From the form (4.35) it is easy to see that in the vicinity of this point we have, to the leading order,
$\Upsilon\approx\left(\frac{3}{2}r\right)^{1/3}\,,\quad\theta\approx-\alpha\,,\quad\Phi_{\text{inst}}\approx\left(\frac{3}{2}\right)^{1/3}\frac{z-it}{(z^{2}+t^{2})^{1/3}}\,$, (4.36)
where $\Upsilon,\theta$ is the parametrization (4.23), and $r,\alpha$ are the polar coordinates in the $(z,t)$ Euclidean plane (4.33). Plugging this into the action in the first line of Eq. (4.25), we find that the would-be divergence of the action integral
@@ 4.5.2 Modes localized on the instanton
The fermionic zero modes can also be counted with the help of the index theorems for the Dirac operator [48, 49, 50]. These index theorems demand, among other things, that the fermion mass (or the coefficient in front of the $\psi\psi$ or $\bar{\psi}\psi$ term) is non-vanishing as long as we are sufficiently far from the vortex location. This is satisfied in the present case, because the fermion mass winds in the complex plane on the kink’s background (4.14b), staying away from zero. However, in the model from Sec. 3, the fermion mass actually vanishes at one point on the kink, arbitrarily far from the instanton-vortex, see footnote 4 on p. 4.
@@ Appendix C Note on a would-be enhanced supersymmetry
Thus, at $\kappa=2$, the bosonic part of the MSTB model coincides with the bosonic part of an $\mathcal{N}=2$ theory. Fermionic part is, of course, different, as the superpotential for the MSTB model (3.8) is still non-holomorphic for any value of $\kappa$.

### v3

@@ 2.2.2 Anomaly in the superpotential
The anomaly in ${\mathcal{N}}=(1,1)$ theories was first discovered in a model with a single real boson field (and its fermion superpartner) and the sine-Gordon model almost 30 years ago. Depending on a specific regularization method used, computations can yield the kink mass which may or may not violate the BPS bound, see e.g. [34, 35, 36]. This issue was later clarified in [30], where supersymmetry-preserving regularization was used.
Here we will outline the derivation of [30] assuming that the number of fields is arbitrary. We start from the Lagrangian (2.1) implying the classical expression for the supercurrent quoted in (2.2). Classically,
@@
due to the fact that $\gamma_{\mu}\gamma^{\nu}\gamma^{\mu}=0$ in two dimensions. As we will see, the right-hand side of (2.15) changes upon the UV regularization. The simplest one is the dimensional regularization [^2] Additional verification is provided by a regularization through higher derivatives [30] and dimensional reduction [37]. to be used below in which $\gamma_{\mu}\gamma^{\nu}\gamma^{\mu}=D-2$.
To carry out the one-loop calculation [^3] Why one loop is sufficient will be explained below., we apply the background field technique; namely we substitute the field $\phi^{a}$ by its background and quantum parts, $\phi^{a}$ and $\chi^{a}$, respectively,
@@ 3.2 $\mathcal{N}=(1,1)$ supersymmetrization
Figure 1: (Super)Potentials for the MSTB model for $\lambda=m=1$. (1(a)) Potential (3.2) at $\kappa=0.8$. (1(b)) and (1(c)) Superpotential (3.8) at $\kappa=0.3$ (the same plot from two different points of view). The values of $\kappa$ are chosen so that the characteristic features are most prominent visually. The thin gray contour lines mark the height of $U$ or $\mathcal{W}$. The profiles of the BPS degenerate kinks (3.15) are shown by the thick solid lines, magenta and green respectively. The sphaleron profile (3.32) is shown by the thick dashed red line. On (1(b)) and (1(c)), the BPS (and semi-BPS) profiles — the thick lines — can be seen as trajectories of viscous honey trickling down on the superpotential landscape (different trajectories correspond to different infinitesimal initial speeds).
@@ 3.5.2 Scalar VEVs and the fate of the ground states
Now let us consider some implications of the correction to the superpotential, Eq. (3.47). We will focus on the case $|\kappa|\leqslant 1$ relevant for the study of degenerate kinks. As in Eq. (2.26), we introduce the full superpotential:
@@
The last term here is the result of the quantum correction (3.47). Note that it is strictly positive, and for large enough $\lambda$ this equation might not have real solutions.
@@
$\lambda_{\text{crit,2}}=\frac{m}{3}\sqrt{\frac{\pi}{2}\left(2+\kappa^{2}+\kappa\sqrt{3(4-\kappa^{2})}\right)}\,.$ (3.56)
@@ 3.6 Instanton (bosonic sector)
Now it is time to discuss the instanton configuration that interpolates between the two degenerate kinks discussed above. This study was initiated in [17]. For now we will focus only on the bosonic sector; the fermionic sector will be discussed later.
@@
Physically, one way to think about an instanton is to consider it a classical trajectory (or a field profile) that interpolates between different ground states of the theory. At the same time, in the model at hand we wish to consider an instanton that interpolates between two kinks. To consolidate these notions, we can put our theory in boundary conditions with a spatial twist with respect to the $\mathbb{Z}_{2}$ transformation $S:\ \phi_{1}\to-\phi_{1}$, see Eq. (3.3a). Recall that this $S$ is a valid symmetry of the bosonic sector of the theory; the fermionic sector will be discussed below. Such a twist puts the theory in the non-trivial topological sector; in particular, the ground states of this theory necessarily have non-trivial topology as well. Since the two degenerate kinks are the global minima of the action with the given boundary conditions (they saturate the Bogomolny bound), we conclude that, in fact, these two kinks are the ground states in this sector.
@@
3.6.1 Bosonic worldline effective action at $\kappa\sim\pm 1$
@@
With the ansatz (3.64), it is straightforward to obtain the effective action for the $\beta(\tau)$. To this end we plug the ansatz (3.64) into the Euclidean action (3.63), expand in the powers of $\beta$, and keep only a few relevant lowest-order terms. Recall that we assume $|\beta(\tau)|<\alpha\ll 1$. Integrating over $z$, we obtain [^9] See [29] for details. the effective quantum mechanical action for $\beta(\tau)$ (Euclidean notation):
$S_{\beta}^{\text{eff}}=\int d\tau\left\{\frac{m}{\lambda^{2}}\frac{1}{2}\dot{\beta}^{2}+V^{\text{eff}}(\beta)\right\}\,$, (3.65)
@@
Therefore, when $\alpha$ is not small, interference effects can become noticeable. In particular, there may be a certain value of the parameter $\alpha$, at which two different one-instanton amplitudes cancel because of the non-trivial relative phase.
@@ 3.6.2 Instanton in the limit $\kappa\to 0$
At the point $\kappa=0$, the kinks disappear, but the instanton solution is formally there — it becomes the global vortex. Its action diverges logarithmically in this limit,
@@ 3.7.1 Overall translational modes of the kink-instanton system
We expect that the kink zero modes from Sec. 3.3.3 smoothly continue through the instanton. This is an overall fermionic mode of the configuration from Fig. 7. When we switch on instantons, both in plain and topologically nontrivial backgrounds, we must perform the Wick rotation, i.e. pass to Euclidean time. Then one fermion zero mode on the kink we discussed becomes a zero mode on the domain line in the plane $\{\tau,z\}$ ($\tau$ is the Euclidean time). Its profile function depends on $z$ but is $t$-independent. The instanton deforms this line locally, near its center $z_{0}$ and $\tau_{0}$, but the boundary behavior of the background field remains the same, which means that the profile is deformed around $z_{0},\tau_{0}$ but the mode survives.
@@ 3.7.2 Modes localized on the instanton
Unfortunately, index theorems cannot be directly applied in the present case, because one mass eigenvalue vanishes at a line in the field space, see footnote 5 on p. 5. This does not allow us to rigorously prove our statement (unlike the $\mathcal{N}=2$ case treated in Sec. 4.5.2). Still, below in Sec. 3.7.4 we present an argument, which, although conjectural, makes a plausible case. But before doing that, we need to recall some facts about $\mathcal{N}=2$ 2d Wess-Zumino models with holomorphic superpotentials.
@@ 4.1 Setup
Here, the prime represents a derivative w.r.t. $\Phi$. Perturbative mass near each of the vacua (4.4) as calculated form the scalar potential $V$ in Eq. (4.3):
$(m_{\text{pt}})^{2}=\frac{4b^{3}}{a}\,.$ (4.5)
@@
The superpotential (4.1) is motivated by $\mathcal{N}=1$ SQCD in 4D, with gauge group $SU(2)$ and one quark flavor. There, $\Phi$ plays the role of the meson superfield. In that context, $b=m$ is the quark mass, and $a=\Lambda^{5}$ in terms of the strong coupling scale. Generalization to the gauge group $SU(N)$ yields the superpotential
@@ 4.2 Two degenerate kinks
$\partial_{z}\theta=-2\frac{b^{3/2}}{a^{1/2}}\sin\theta\,.$ (4.13)
@@
$\theta=\pm 2\;\mathrm{arctan}(e^{-m_{\text{pt}}\,(z-z_{0})})\,$, (4.14)
@@
$\Phi_{\text{u-kink}}(z)=\Phi_{+}\cdot\frac{e^{m_{\text{pt}}\,z}+i}{e^{m_{\text{pt}}\,z}-i}\,$, (4.15a)
$\Phi_{\text{d-kink}}(z)=\Phi_{+}\cdot\frac{e^{m_{\text{pt}}\,z}-i}{e^{m_{\text{pt}}\,z}+i}\,.$ (4.15b)
Here, $m_{\text{pt}}$ is naturally the mass of perturbative excitations near each of the vacua, see Eq. (4.5). In terms of the real component fields (4.2), the two kinks become
$\phi_{1}(z)=\Phi_{+}\tanh(m_{\text{pt}}\,z)\,,\quad\phi_{2}(z)=\pm\frac{\Phi_{+}}{\cosh(m_{\text{pt}}\,z)}\,$, (4.16)
cf. Eq. (3.15). The plus/minus sign in Eq. (4.16) corresponds to the u/d-kink respectively.
Both kinks in Eq. (4.15b) interpolate from $\Phi_{-}$ at $z\to$-$\infty$ to $\Phi_{+}$ at $z\to+\infty$, see Fig. 9. The first kink $\Phi=\Phi_{\text{u-kink}}(z)$ goes through the upper half-plane on the field space; the phase of $\Phi$ (i.e. $\theta$) interpolates from $\pi$ to 0 clockwise. The second kink $\Phi=\Phi_{\text{d-kink}}(z)$ goes through the lower half-plane on the field space; the phase of $\Phi$ interpolates from $\pi$ to 0 going counter-clockwise.
@@ 4.3.1 First order equations for sphaleron profile
$z-z_{0}=\frac{1}{b}\left[\Phi_{+}\,\text{atanh}\frac{\Phi_{\text{sph}}(z)}{\Phi_{+}}-\Phi_{\text{sph}}(z)\right]\,$, (4.17)
where the vacuum value $\Phi_{+}$ is defined in (4.4), and $z_{0}$ is an arbitrary constant (the sphaleron center). For a visual representation, see Fig. 9. Note that, contrary to a kink, the sphaleron (4.17) satisfies Eq. (4.10) with the choice of sign as “-”.
@@
$\Phi_{\text{sph}}(z)\approx(3a(z-z_{0}))^{1/3}\,,\quad\Phi_{\text{sph}}^{\prime}(z)\sim\frac{1}{(z-z_{0})^{2/3}}\,.$ (4.18)
@@ 4.3.2 Sphaleron energy and BPS bound
$\delta\Phi(z)_{\text{transl}}\sim\partial_{z}\Phi_{\text{sph}}(z)\,.$ (4.21)
@@ 4.4.1 Topological charge
Here, $M_{\text{kink}}$ is given by the absolute value of the central charge (4.12). The term $\int d\tau M_{\text{kink}}$ in (4.25) comes from $2\int dz\,d\tau\,[\partial_{\zeta}\mathcal{W}+c.c.]$. Formally, this term is divergent, as we need to integrate the kink mass over all of time. However, with given boundary conditions at $z\to\pm\infty$, a kink represents a ground state, so we subtract that contribution away, as in Sec. 3.
The term $S_{\text{inst}}$ in (4.25) gives the instanton action; it is given by
@@ 4.4.2 BPS profile equations and the singularity
Note that the degenerate kinks from Eq. (4.15b) also satisfy these equations (assume $\tau$-independent fields and compare Eq. (4.32) with the kink equation (4.10)).
Let us introduce polar coordinates on the Euclidean plane $(z,\tau)$ according to
$z=r\cos\alpha\,,\quad\tau=r\sin\alpha\,.$ (4.33)
@@
Previously we saw that the sphaleron had a singularity, which led to a divergence of the sphaleron’s energy. One might worry, what is the fate of the instanton, since the instanton configuration inevitably passes through the point $\Phi=0$ in the field space; let’s say that this happens at the origin of the $(z,\tau)$ plane. From the form (4.35) it is easy to see that in the vicinity of this point we have, to the leading order,
$\Upsilon\approx\left(\frac{3}{2}ar\right)^{1/3}\,,\quad\theta\approx-\alpha\,,\quad\Phi_{\text{inst}}\approx\left(\frac{3}{2}a\right)^{1/3}\frac{z-i\tau}{(z^{2}+\tau^{2})^{1/3}}\,$, (4.36)
where $\Upsilon,\theta$ is the parametrization (4.23), and $r,\alpha$ are the polar coordinates in the $(z,\tau)$ Euclidean plane (4.33). Plugging this into the action in the first line of Eq. (4.25), we find that the would-be divergence of the action integral
@@ 4.5.2 Modes localized on the instanton
The fermionic zero modes can also be counted with the help of the index theorems for the Dirac operator [53, 54, 55]. These index theorems demand, among other things, that the fermion mass (or the coefficient in front of the $\psi\psi$ or $\bar{\psi}\psi$ term) is non-vanishing as long as we are sufficiently far from the vortex location. This is satisfied in the present case, because the fermion mass winds in the complex plane on the kink’s background (4.15b), staying away from zero. However, in the model from Sec. 3, the fermion mass actually vanishes at one point on the kink, arbitrarily far from the instanton-vortex, see footnote 5 on p. 5.
@@ Appendix C Note on a would-be enhanced supersymmetry
Thus, at $\kappa=2$, the bosonic part of the MSTB model coincides with the bosonic part of an $\mathcal{N}=2$ theory. The fermionic part is, of course, different, as the superpotential for the MSTB model (3.8) is still non-holomorphic for any value of $\kappa$.